a particle is thrown vertically upwars. its velocity at half of the height is 10m/s.what is the mximum height attained by it ?
1
Expert's answer
2012-10-10T11:17:29-0400
Let initial velocity is U and maximum height is h.
At maximum height velocity will be zero:
0 =U2 -2gh
U2 =2gh
Now given velocity is = 10m/sec at height h/2
so 102 = U2 -2gh/2
100 =2gh -gh
h = 100/10
h =10 m